### v1

@@ Abstract
Current causal discovery approaches require restrictive model assumptions or assume access to interventional data to ensure structure identifiability. These assumptions often do not hold in real-world applications leading to a loss of guarantees and poor accuracy in practice. Recent work has shown that, in the bivariate case, Bayesian model selection can greatly improve accuracy by exchanging restrictive modelling for more flexible assumptions, at the cost of a small probability of error. We extend the Bayesian model selection approach to the important multivariate setting by making the large discrete selection problem scalable through a continuous relaxation. We demonstrate how for our choice of Bayesian non-parametric model, the Causal Gaussian Process Conditional Density Estimator (CGP-CDE), an adjacency matrix can be constructed from the model hyperparameters. This adjacency matrix is then optimised using the marginal likelihood and an acyclicity regulariser, outputting the maximum a posteriori causal graph. We demonstrate the competitiveness of our approach on both synthetic and real-world datasets, showing it is possible to perform multivariate causal discovery without infeasible assumptions using Bayesian model selection.
@@ 1 Introduction
In many systems, for example, protein signalling networks, variables causally relate to each other as changing a variable only modifies certain variables (Sachs et al., 2005). Uncovering these unique underlying causal structures from data can allow us to gain new insights in a wide range of fields, from Biology (Sachs et al., 2005) to Medicine (Feuerriegel et al., 2024) to Economics (Hicks et al., 1980). There are two main approaches to learning causal structures from data. The first approach assumes a restricted model class for the data generating process (Peters et al., 2017, Ch. 4), but its guarantees often fail when these assumptions don’t hold. The second approach requires interventional data for all variables, which can be costly, ethically challenging, or even impossible (Li et al., 2019). Consequently, learning a unique causal structure often relies on impractical assumptions, limiting its real-world applicability.
Recent work has shown that Bayesian model selection can be used to learn a unique causal structure without interventional data while broadening the flexibility of usable model classes (Dhir et al., 2024). This comes at a cost of loosened identifiability guarantees but with the ability to posit more realistic assumptions. In fig. 1, we illustrate a causal relationship in the Sachs dataset (Sachs et al., 2005) that our less restrictive model captures accurately, while most methods assuming additive noise (ANM) fail (Hoyer et al., 2008). Figure 1(a) demonstrates that the ANM assumption is inappropriate, causing its guarantees to break down. In contrast, Bayesian model selection enables the use of models with more suitable assumptions (fig. 1(b)). However, the current formulation of Bayesian model selection for causal discovery requires computing and comparing the posteriors of all possible causal graphs. As the number of graphs grows super-exponentially with the number of variables, this approach becomes computationally intractable for large graphs.
Figure 1: The figure shows model samples for the PKA → ERK pathway from the Sachs dataset (Sachs et al., 2005). (a) Samples from a Gaussian Process (GP) with the ANM assumption. (b) Samples from the GP-CDE model. The ANM assumption is unsuitable here as it cannot capture the varying noise, whereas the GP-CDE model assumptions can.
A potential solution lies in continuous optimisation causal discovery approaches (Zheng et al., 2018). These approaches interpret causal discovery as a single optimisation problem that is amenable to gradient based methods. So far, methods making use of continuous optimisation have been limited to using restricted model classes or interventional data. In this work, we propose a scheme to extend Bayesian model selection based causal discovery to multiple variable systems using continuous optimisation — facilitating more flexible causal discovery at larger scales.
We start with a Bayesian non-parametric model that can model non-Gaussian and heteroscedastic densities, the Gaussian Process conditional density estimator (GP-CDE) (Titsias and Lawrence, 2010; Lalchand et al., 2022). We show that certain hyperparameters of this model can be interpreted as an adjacency matrix for a graph, allowing for continuously parametrising the space of graphs. These hyperparameters can then be optimised over, using a loss derived from Bayesian model selection, and a regulariser to ensure the final adjacency represents a valid acyclic graph. Priors placed over these hyperparameters can also be interpreted as priors over graphs, leading to a scheme that returns the maximum a-posteriori (MAP) causal graph. We show the practical advantage of this method with competitive performance across synthetic and real datasets. Our work shows that it is beneficial in practice to give up on strict identifiability results under unrealistic assumptions for realistic assumptions, even if these imply a small but irreducible probability of error.
@@ 2 Background
Our work is concerned with recovering causal structure from data. Here, we outline material relevant to this task and the assumptions required.
@@ 2.1 Causal Model
We make the common assumption that there are no hidden confounders. We assume that data $\mathbf{X}\in\mathbb{R}^{N\times D}$ can be explained by a Causal Model, which we refer to as $\mathcal{M}_{\mathcal{G}}$. A causal model can be formally defined by a directed acyclic graph (DAG) with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$, denoted $\mathcal{G}=(\mathcal{V},\mathcal{E})$, along with a set of conditional distributions $\mathcal{C}:=\prod_{i\in\mathcal{V}}\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$, where $\mathrm{PA}_{\mathcal{G}}(i)$ denotes the parents of node $i$ in $\mathcal{G}$, and $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ is the set of conditionals for variable $i$. An element of $\mathcal{C}$ induces a joint distribution over $D$ random variables $X=\{X_{1},\ldots,X_{D}\}$ when its constituent members are multiplied together. That is, for $(P_{i}:i\in\mathcal{V})\in\mathcal{C}$ the joint factorises as $\prod_{i=1}^{D}P_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{G}}(i)})$. We denote the set of all induced joints as =$\mathcal{F}=\{\prod_{i=1}^{D}P_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{G}}(i)}):(P_{i}:i\in\mathcal{V})\in\mathcal{C}\}$ (Dhir et al., 2024). We also assume likelihood modularity (Geiger and Heckerman, 2002), that is for variable $i$ if $\mathrm{PA}_{\mathcal{G}}(i)=\mathrm{PA}_{\mathcal{G}^{\prime}}(i)$ for $\mathcal{G}\neq\mathcal{G}^{\prime}$, then $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}=\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}^{\prime}}(i)}$. This means that for graphs where variables have the same parents, we assume the same set of conditional distributions.
The causal model is assumed to satisfy the Independent Causal Mechanism (ICM) assumption (Janzing and Schölkopf, 2010). This states that the terms in the product $\prod_{i=1}^{D}P_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{G}}(i)})$, that constitutes the factorisation according to $\mathcal{G}$, are composed of autonomous modules. This implies that changing one term (such as under an intervention) leaves all other terms unchanged. This property does not necessarily hold in any other factorisation of the joint (Peters et al., 2017).
As variables are dependent only on their parents in a graph by construction, a causal model satisfies the Markov assumption. This states that d-separations in $\mathcal{G}$ between variables implies conditional independence between the same variables (Pearl, 2009). We also assume faithfulness: any conditional independences between variables are represented by d-separations in $\mathcal{G}$ (Pearl, 2009). Together, faithfulness and Markov assumptions imply a one-to-one relationship between independences in the distributions and d-separations in the graph (Peters et al., 2017). These assumptions are enough to recover a causal structure from data up to a Markov equivalence class (MEC), that is the class of causal structures that have the same d-separations.
@@ 2.2 Learning Causal Structure
Formally, the task is to recover a DAG $\mathcal{G}$ given a dataset of $N$ samples with $D$ variables, $\mathbf{X}\in\mathbb{R}^{N\times D}$. As stated in section 2.1, without any interventional data and equipped with only the faithfulness and Markov assumptions, only the MEC of a DAG can be recovered (Pearl, 2009). To recover a unique DAG, and differentiate within an MEC, we require additional assumptions.
@@ 2.3 Learning Causal Structure with Functional Restrictions
One way to recover a DAG is to impose prior restrictions on the allowable conditional distributions $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$. These restrictions are chosen so that the joint distributions implied by a causal model cannot be expressed by any other causal model— specifically, for $P\in\mathcal{F}_{\mathcal{G}}$ it holds that $P\not\in\mathcal{F}_{\mathcal{G}^{\prime}}$ for any $\mathcal{G}^{\prime}\neq\mathcal{G}$ (Guyon et al., 2019, Ch. 2). Thus if data is sampled from one of the causal models, the causal structure can be identified with probability 1. Zhang et al. (2015) show that the maximum likelihood score is sufficient to identify the causal structure with these restricted models. As an example, additive noise models (ANM) (Hoyer et al., 2008) restrict all $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ to the form $f(X_{\mathrm{PA}_{\mathcal{G}}(i)})+\epsilon$ for some arbitrary non-linear function $f$, and arbitrarily distributed $\epsilon$. If we consider a different graph $\mathcal{G}^{\prime}$ but same restrictions on $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}^{\prime}}(i)}$, it is not possible to approximate the joint induced by the original set of conditional distributions $\prod_{i\in\mathcal{V}}\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$.
@@ 2.4 Learning Causal Structure with Bayesian Model Selection
Bayesian model selection has previously been used with linear models to learn an MEC (Heckerman, 1995; Heckerman and Geiger, 1995; Heckerman et al., 1995; Geiger and Heckerman, 2002). Following other attempts (Friedman and Nachman, 2000; Stegle et al., 2010), Dhir et al. (2024) showed that Bayesian model selection can allow for distinguishing between distinct casual models even if they can express the same distribution, such as when using non-parametric model classes. Here, the strict identifiability of restricted model classes no longer holds, although the probability of error may be small depending on the model used and can be estimated empirically (Dhir et al., 2024) (discussed further in appendix A). The advantage of this approach is that it allows for using more flexible model classes to learn causal structures, potentially mitigating the misspecification issues discussed in section 2.3, where the guarantees of restricted model classes do not hold. Our work takes advantage of these insights.
@@
where $P(\mathcal{M}_{\mathcal{G}})$ is the prior over the causal model. To indicate no preference over causal direction between variables, the prior is chosen to be uninformative about graphs in the same MEC. The term $P(X|\mathcal{M}_{\mathcal{G}})$ is known as the marginal likelihood. We select the most likely causal model $\mathcal{M}_{\mathcal{G}}^{*}$ based on the highest posterior value (Kass and Raftery, 1995)
@@
To calculate the marginal likelihood, a prior (denoted $\pi$) over the distributions in the causal model must be defined. The only constraint placed on the prior is that it should encode the ICM assumption (section 2.1)(Stegle et al., 2010; Dhir et al., 2024). The ICM assumption implies that knowledge about any distribution $P_{i}\in\mathcal{C}_{i|\mathrm{PA}(i)}$ should not inform any other $P_{j}\in\mathcal{C}_{j|\mathrm{PA}(j)}$ for $j\neq i$. This is achieved by defining factorised priors on each variable distribution set — $\pi_{i}\in\mathcal{P}(\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)})$ where $\pi_{i}$ is the prior on the distributions for variable $i$, and $\mathcal{P}$ is the set of all distributions over an object. Given the prior, the marginal likelihood is
@@
The marginal likelihood score can be interpreted as how well a prior, and the encoded ICM assumption within it, explains a given dataset (Dhir et al., 2024). The prior indicates the weight given to distributions in $\mathcal{C}_{i|\mathrm{PA}{\mathcal{G}}(i)}$ used to explain the data, effectively softening the hard restrictions discussed in section 2.3. As the prior is a distribution, it must also sum to one which means that, unlike maximum likelihood, it cannot equally weigh every element in $\mathcal{C}_{i|\mathrm{PA}{\mathcal{G}}(i)}$. This means that although multiple causal models can represent a given distribution, the prior and the encoded ICM assumption favour certain models over others—specifically, those whose prior and encoded ICM assumption align with that of the observed data.
Dhir et al. (2024) only consider Bayesian model selection for the two variable case. For a higher number of variables, the number of DAGs grows super exponentially. This makes the procedure of comparing the marginal likelihood of all the causal models with distinct DAGs prohibitive. In this paper, we investigate the feasibility of scaling this approach for multiple variables.
2.5 Continuous Optimisation for Learning Causal Structure
That is, finding the graph that optimises the score under the constraint that the graph must be a valid DAG. This converts causal structure learning into a single optimisation problem, tackling the scalability issue. However, ensuring that the graph is acyclic is nontrivial with gradient based methods.
Zheng et al. (2018) solve this by first encoding a graph with a weighted adjacency matrix $\mathbf{A}\in\mathbb{R}^{D\times D}_{\geq 0}$, then a measure of acyclicity is introduced
$h(\mathbf{A})=\text{Tr}(e^{\mathbf{A}})-D,$ (5)
where $e^{\mathbf{A}}$ is the matrix exponential. The function $h:\mathbf{A}\to\mathbb{R}_{\geq 0}$ is a measure of the number of weighted directed walks that allow for returning to a starting node. Hence, $h(\mathbf{A})=0$ implies that there are no directed cycles in the adjacency $\mathbf{A}$ and that the graph implied by $\mathbf{A}$ is a valid DAG.
Equation 4 can now be written as an optimisation problem with the differentiable constraint $h(\mathbf{A})=0$. Zheng et al. (2018) use the augmented Lagrangian method to convert the above into a sequence of unconstrained optimisation problems (Nemirovsky, 1999). Solutions to the sequence of subproblems are solutions of the original constrained optimisation problem. Each subproblem is
$\operatorname*{\arg\!\max}_{\mathbf{A}}\mathcal{S}(\mathcal{G}^{\mathbf{A}})-\frac{\rho_{t}}{2}|h(\mathbf{A})|^{2}-\alpha_{t}h(\mathbf{A}),$ (6)
where $t$ is the problem index, $\rho_{t}$ and $\alpha_{t}$ are coefficients of the method, while $\mathcal{G}^{\mathbf{A}}$ denotes a DAG $\mathcal{G}$ with adjacency matrix $\mathbf{A}$. After each subproblem is solved, the coefficients are increased (see section B.3). A solution $\mathbf{A}^{*}$ is found when $h(\mathbf{A}^{*})<\epsilon$ for some convergence threshold $\epsilon>0$.

### v2

@@ Abstract
Current causal discovery approaches require restrictive model assumptions in the absence of interventional data to ensure structure identifiability. These assumptions often do not hold in real-world applications leading to a loss of guarantees and poor performance in practice. Recent work has shown that, in the bivariate case, Bayesian model selection can greatly improve performance by exchanging restrictive modelling for more flexible assumptions, at the cost of a small probability of making an error. Our work shows that this approach is useful in the important multivariate case as well. We propose a scalable algorithm leveraging a continuous relaxation of the discrete model selection problem. Specifically, we employ the Causal Gaussian Process Conditional Density Estimator (CGP-CDE) as a Bayesian non-parametric model, using its hyperparameters to construct an adjacency matrix. This matrix is then optimised using the marginal likelihood and an acyclicity regulariser, giving the maximum a posteriori causal graph. We demonstrate the competitiveness of our approach, showing it is advantageous to perform multivariate causal discovery without infeasible assumptions using Bayesian model selection.
Machine Learning, ICML
@@ 1 Introduction
In many systems, such as protein signalling networks, variables causally relate to each other, since changing a variable only modifies certain variables (Sachs et al., 2005). Uncovering these unique underlying causal structures from data can allow us to gain new insights in a wide range of fields, from Biology (Sachs et al., 2005) to Medicine (Feuerriegel et al., 2024) to Economics (Hicks et al., 1980). There are two main approaches for learning the ground truth causal structure from data. The first approach assumes a restricted model class for the data generating process (Peters et al., 2017, Ch. 4), but its guarantees often fail when these assumptions do not hold. The second approach requires interventional data for all variables, which can be costly, ethically challenging, or even impossible (Li et al., 2019). Consequently, learning a unique causal structure often relies on impractical assumptions, limiting its real-world applicability.
Recent work has shown that, in the absence of interventional data and in the bivariate case, Bayesian model selection can more accurately infer the unique causal structure of a system (Dhir et al., 2024). Unlike previous methods, including other Bayesian methods [^1] These methods are usually concerned with inferring a distribution instead of a single causal structure., which impose constraints on the function class, this framework allows the use of more flexible model classes. This comes at a cost of loosened identifiability guarantees but with the ability to posit more realistic assumptions. However, Dhir et al. (2024) only study the bivariate case, while most applications require understanding causal relationships between multiple variables. It is thus an open question whether the advantages of this method hold in the multivariate case. Furthermore, the approach in Dhir et al. (2024) requires computing and comparing the posteriors of all possible causal graphs. As the number of graphs grows super-exponentially with the number of variables, this becomes computationally intractable for scalability.
A potential solution lies in continuous optimisation causal discovery approaches (Zheng et al., 2018; Nazaret et al., 2024). These approaches interpret causal discovery as a single continuous optimisation problem that is amenable to gradient based methods. The emphasis of this line of work has been either on scaling to a larger number of variables or on loosening model assumptions. Our contribution tackles the latter. We argue that the potential advantages of Bayesian methods enable them to outperform their non-Bayesian counterparts.
We investigate whether and how well Bayesian model selection can recover the ground truth causal structure in multivariable systems. We use a model based on the Gaussian Process conditional density estimator (GP-CDE) (Titsias & Lawrence, 2010; Lalchand et al., 2022) that is more flexible than traditional identifiable models that rely on restrictive model assumptions. Interpreting the model hyperparameters as a graph, we optimise over them using a Bayesian model selection-based loss and a regulariser to enforce acyclicity, ultimately producing the maximum a posteriori (MAP) causal graph. Our experiments address two critical questions: (1) How much performance is lost compared to enumerating every causal graph? (2) How does our scalable Bayesian model selection compare to existing multivariate causal discovery methods in terms of performance? Our results show that Bayesian model selection outperforms methods enforcing strict identifiability even at larger scales. While it introduces a small error probability, the flexibility afforded enables more realistic and practical multivariable causal discovery, advancing toward real-world applications.
@@ 2 Background
Here, we outline the assumptions relevant to the task of causal discovery. Additional background is in Appendix A.
@@ 2.1 Causal Model
We assume that data $\mathbf{X}\in\mathbb{R}^{N\times D}$ can be explained by a Causal Model, which we refer to as $\mathcal{M}_{\mathcal{G}}$. We make the common assumption that there are no hidden confounders.
Definition 2.1.
(Consistent with Pearl (2009) but using notation from Dhir et al. (2024)) A Causal Model $\mathcal{M}_{\mathcal{G}}:=(\mathcal{G},\mathcal{C}_{\mathcal{G}})$ is defined as a tuple containing a DAG $\mathcal{G}$ with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$, along with a set of conditional distributions for each variable $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$, where $\mathrm{PA}_{\mathcal{G}}(i)$ denotes the parent index set of node index $i$ in $\mathcal{G}$. The set of all possible conditionals is the Cartesian product $\mathcal{C}_{\mathcal{G}}:=\prod_{i\in\mathcal{V}}\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$. The set of all possible conditional distributions induces a set of joint distributions denoted $\mathcal{F}_{\mathcal{G}}$.
An element of $\mathcal{C}_{\mathcal{G}}$ induces a joint distribution over $D$ random variables $X=\{X_{1},\ldots,X_{D}\}$ when its constituent members are multiplied together. That is, for an element $(P_{i}:i\in\mathcal{V})\in\mathcal{C}_{\mathcal{G}}$ the joint is $\prod_{i=1}^{D}P_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{G}}(i)})$. Hence $\mathcal{C}_{\mathcal{G}}$ induces a set of joint distributions $\mathcal{F}_{\mathcal{G}}=\{\prod_{i=1}^{D}P_{i}(X_{i}|X_{\mathrm{PA}_{\mathcal{G}}(i)}):(P_{i}:i\in\mathcal{V})\in\mathcal{C}_{\mathcal{G}}\}$ (Dhir et al., 2024). We also assume likelihood modularity (Geiger & Heckerman, 2002), that is for variable $X_{i}$ if $\mathrm{PA}_{\mathcal{G}}(i)=\mathrm{PA}_{\mathcal{G}^{\prime}}(i)$ for $\mathcal{G}\neq\mathcal{G}^{\prime}$, then $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}=\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}^{\prime}}(i)}$. This means that for graphs where variables have the same parents, we assume the same set of conditional distributions.
@@ 2.2 Learning Causal Structure
Formally, the task is to recover a DAG $\mathcal{G}$ given a dataset of $N$ samples with $D$ variables, $\mathbf{X}\in\mathbb{R}^{N\times D}$. Without any interventional data, only the Markov equivalence class (MEC) of a DAG can be recovered with the faithfulness assumption (Pearl, 2009) and with infinite data (Appendix A). To recover a unique DAG, and differentiate within an MEC, we require additional assumptions.
@@ 2.3 Learning Causal Structure with Functional Restrictions
One way to recover a DAG is to impose a priori restrictions on the allowable conditional distributions $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$. These restrictions are chosen so that the joint distributions implied by a causal model cannot be expressed by any other causal model— specifically, for $P\in\mathcal{F}_{\mathcal{G}}$ it holds that $P\not\in\mathcal{F}_{\mathcal{H}}$ for any $\mathcal{H}\neq\mathcal{G}$ (Guyon et al., 2019, Ch. 2). Thus, if data is sampled from one of the causal models, the causal structure can be identified. As an example, additive noise models (ANM) (Hoyer et al., 2008) restrict all $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ to the form $f(X_{\mathrm{PA}_{\mathcal{G}}(i)})+\epsilon$ for some arbitrary non-linear function $f$, and arbitrarily distributed $\epsilon$. If we consider a different graph $\mathcal{H}$ but the same ANM restrictions on $\mathcal{C}_{i|\mathrm{PA}_{\mathcal{H}}(i)}$, it is not possible to approximate the joint induced by the original set of conditional distributions $\mathcal{C}_{\mathcal{G}}=\prod_{i\in\mathcal{V}}\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ using $\mathcal{C}_{\mathcal{H}}=\prod_{i\in\mathcal{V}}\mathcal{C}_{i|\mathrm{PA}_{\mathcal{H}}(i)}$. This allows the graph $\mathcal{G}$ to be identified. For a lot of these restrictions, the maximum likelihood score can identify the causal structure (Zhang et al., 2015; Peters & Bühlmann, 2014; Immer et al., 2023).
@@ 2.4 Learning Causal Structure with Bayesian Model Selection
Restricted function classes are required for identifiability within an MEC as it implies that, with infinite data, only one of the causal models can approximate the data distribution and thus achieve a higher score such as the maximum likelihood. Without restrictions, multiple causal structures can yield the same maximum likelihood score, making it impossible to distinguish between them. However, using such an unrestricted causal model with a Bayesian model selection-based score can restore the ability to differentiate between causal structures. This is because, while multiple structures may share the same maximum likelihood score, they typically yield distinct scores under Bayesian model selection. This insight was demonstrated for the two-variable case in Dhir et al. (2024), building on prior efforts (Friedman & Nachman, 2000; Stegle et al., 2010). While this approach no longer guarantees strict identifiability and admits a small probability of error, the error probability can be empirically estimated and remains low with appropriate models (Dhir et al., 2024). The key advantage of this framework lies in its flexibility. By enabling the use of broader model classes, it mitigates the risk of misspecification, which is a common limitation where the assumptions and guarantees of restricted models fail (Section 2.3). Our work builds on these insights.
@@
where $P(\mathcal{M}_{\mathcal{G}})$ is the prior over the causal model. As we assume no preference over causal direction between variables, our prior belief is equal over graphs in the same MEC. The posterior within an MEC is then dependent only on the term $P(X|\mathcal{M}_{\mathcal{G}})$, known as the marginal likelihood. Given that our task is to infer the single ground truth causal structure, we select the causal model $\mathcal{M}_{\mathcal{G}}^{*}$ based on the highest posterior probability (Kass & Raftery, 1995)
@@
To calculate the marginal likelihood, a prior (denoted $\pi$) over the distributions in the causal model must be specified. This defines a Bayesian Causal Model.
Definition 2.2.
(Dhir et al., 2024) A Bayesian Causal Model, denoted as $(\mathcal{M}_{\mathcal{G}},\pi)$, is defined as a causal model $\mathcal{M}_{\mathcal{G}}$ (Definition 2.1) with a prior distribution $\pi$ over $\mathcal{C}_{\mathcal{G}}$.
Each Bayesian causal model should encode the common and foundational principle that a change in one causal mechanism should not affect any others (Aldrich, 1989; Pearl, 2009). This can be formalised by the independent causal mechanism (ICM) assumption (Janzing & Schölkopf, 2010). The ICM assumption implies that, in the causal factorisation, knowledge about any distribution $P_{i}\in\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)}$ should not inform any other $P_{j}\in\mathcal{C}_{j|\mathrm{PA}_{\mathcal{G}}(j)}$ for $j\neq i$. In the Bayesian approach this implies defining factorised priors on each set of variable distributions — $\pi=\prod_{i\in\mathcal{V}}\pi_{i}$, where $\pi_{i}\in\mathcal{P}(\mathcal{C}_{i|\mathrm{PA}_{\mathcal{G}}(i)})$ is the prior on the distributions for variable $X_{i}$, factorised according to the causal structure $\mathcal{G}$, and $\mathcal{P}$ is the set of all distributions over an object. This is the only constraint we place on the prior.
Multiple Bayesian causal models may achieve the same likelihood score given infinite data, but the prior and the encoded ICM assumption play a key role in yielding distinct marginal likelihood scores. The ICM assumption, which encodes the causal factorisation in the prior, typically does not hold in any other factorisation (Dhir et al., 2024). To see this, consider the model with graph $X\to Y$, with $\mathcal{C}_{X}$ parametrised by $\theta$ and $\mathcal{C}_{Y|X}$ parametrised by $\phi$. Here, placing a prior on $\mathcal{C}_{X\to Y}$ is equivalent to placing a prior on the parameters $\theta$ and $\phi$. ICM implies that the priors over $\theta,\phi$ factorise, which means the joint factorises as
$\pi(\theta)\pi(\phi)P(X|\theta)P(Y|X,\phi).$ (3)
We can find the reverse factorisation of the variables for the same model using Bayes’ rule. Here, in general, it is not possible to parametrise $\mathcal{C}_{Y}$ and $\mathcal{C}_{X|Y}$ with independent parameters (Dhir et al., 2024) [^2] Exceptions exist, notably normalised linear Gaussian models which are known to be unidentifiable.
[…]
$=$ $\pi(\alpha,\beta)P(Y|\alpha)P(X|Y,\beta),$ (5)
where $\alpha,\beta$ are some reparametrisations. Hence, if $\pi(\alpha,\beta)\neq\pi(\alpha)\pi(\beta)$, the model with the graph $X\to Y$ only satisfies the ICM assumption, as specified by factorised priors on the set of conditionals, in the causal factorisation of the joint $\{P_{i}(X)P_{j}(Y|X):P_{i}\in\mathcal{C}_{X},P_{j}\in\mathcal{C}_{Y|X}\}$.
Given a prior, the marginal likelihood is
@@
The marginal likelihood score has an automatic Occam’s razor effect that balances model fit along with complexity (MacKay, 2003, Ch. 28). This means that although multiple causal models may have parameter settings that fit a given distribution, some models may provide a simpler explanation than others. Specifically, those whose prior, and encoded ICM assumption, align with properties of the data generating process. This allows for distinguishing between causal models even with loosening the restrictions discussed in Section 2.3.
Dhir et al. (2024) only consider Bayesian model selection for the two variable case. It is not clear how this approach will perform when the number of variables is increased, especially in contrast with competing methods. We study the multiple variable case where we address the following challenges: First, we use the theory in Dhir et al. (2024) to clearly show that Bayesian model selection can be applied to causal discovery with multiple variables. Secondly, as the number of variables increases, the number of possible DAGs grows super-exponentially, making direct comparison of marginal likelihoods across all DAGs infeasible. To address this, we propose a scalable approach for applying Bayesian model selection to multiple-variable causal discovery.
3 Distinguishability with Multiple Variables
We use theory from Dhir et al. (2024) to show that in the multiple variable case, Bayesian causal models can find the correct DAG using only priors with the ICM condition, rather than hard model restrictions. We proceed by first showing that Bayesian causal models can be identified up to an MEC. Then, we show that Bayesian causal models can differentiate within an MEC, without functional restrictions, but with some probability of error. We only make remarks on key results in this section, but the full set of assumptions, theorems, and proofs is in Appendix B. Note that these definitions hold in the population setting.
First, we define the probability of making an error, given datasets from the chosen Bayesian causal model.
Definition 3.1.
(Dhir et al., 2024) Given a score $\mathcal{S}$, such as the marginal likelihood, we define the probability of error for a model $\mathcal{M}_{\mathcal{G}}$ as
[…]
where =$\mathcal{R}=\{X:\mathcal{S}(X|\mathcal{M}_{\mathcal{H}})>\mathcal{S}(X|\mathcal{M}_{\mathcal{G}})\text{ for }\mathcal{H}\neq\mathcal{G}\}$, and $\mathcal{S}(X|\mathcal{M}_{\mathcal{G}})$ is the score achieved by model $\mathcal{M}_{\mathcal{G}}$ on dataset $X$.
That is, given datasets $X$ from a Bayesian causal model with DAG $\mathcal{G}$, the integral over datasets where the wrong causal model is chosen by the score. For Bayesian causal models, this quantifies the overlap in the posteriors. Note that, similar to previous work, this definition assumes that the data is sampled from one of the models under consideration. The probability of error can thus be estimated empirically for Bayesian causal models by sampling a graph, then sampling datasets from a chosen Bayesian causal model, and inferring the causal graph using the chosen score (Dhir et al., 2024). An identifiable model implies a probability of error of zero in the infinite data limit (Dhir et al., 2024; Guyon et al., 2019). In contrast, an unidentifiable model will have a probability of error equal to that of a uniformly random chosen graph. Similarly, we say a model is distinguishable if the probability of error is less than that of a uniformly chosen random graph.
Remark 3.2.
With the faithfulness assumption, Bayesian model selection can identify a Bayesian causal model up to the MEC of its graph (Theorem B.3).
Although this is well known for certain parametric models (Heckerman, 1995; Chickering, 2002), we show that it is true for the general non-parametric case.
The advantage of the Bayesian approach is apparent when considering cases where the underlying causal model is unidentifiable with maximum likelihood. This is the case for graphs within an MEC, where functional restrictions have not been made.
Remark 3.3.
Suppose that the ICM assumption only holds in the causal factorisation of the Bayesian causal model. Bayesian model selection can then distinguish within an MEC (Theorem B.7).
Thus, without restrictions and if the ICM condition only holds in the causal factorisation, Bayesian model selection can be used to distinguish Bayesian causal models, in that the probability of error is lower than a uniformly random chosen graph. The model class we choose only satisfies the ICM condition in the causal factorisation (Dhir et al., 2024, Appendix D.3). The probability of error for our chosen model class was shown to be very low for the two variable case in Dhir et al. (2024, Section 4.3). We estimate the probability of error for the multivariate case and show that it is low in Section 6.1.
As is common with identifiability results, the probability of error estimate assumes that the data is generated from one of the candidate causal models. However, this assumption may not hold in practice. Dhir et al. (2024, Section 4.4) demonstrate that the model’s estimated probability of error and the true probability of error (from the actual data-generating process) are bounded by the total variation distance between the model and the true distribution. Consequently, mild violations of model assumptions, such as the prior, can still yield accurate error probability estimates.